# Relative Obstructions and Spectral Diagnostics for Sheaves on Cell Complexes

### v1

@@ 4.2.3 Equivalence under compatibility
Let $K$ be a finite simplicial complex and let $\varepsilon:\mathcal{F}\to\mathcal{W}$ be a cellular sheaf morphism, by assuming the compatibility condition 4.2.1, and write $\epsilon^{\ast}:C^{\ast}(K;\mathcal{F})\to C^{\ast}(K;\mathcal{W})$ for the induced cochain map. Let $\widehat{K}$ denote the geometric cone on $K$ with apex vertex $\ast$, and let $\widehat{\mathcal{F}}$ be the cone sheaf on $\widehat{K}$ defined by \$4.2.2.
@@
$\Phi:C^{n}(\widehat{K};\widehat{\mathcal{F}})\cong\Bigl(\bigoplus_{\tau\in K_{n}}\mathcal{F}(\tau)\Bigr)\ \oplus\ \Bigl(\bigoplus_{\sigma\in K_{n-1}}\mathcal{W}(\sigma)\Bigr)\cong C^{n}(K;\mathcal{F})\ \oplus\ C^{n-1}(K;\mathcal{W})=\mathrm{Cone}(\epsilon^{\ast})[1]^{n},$
where $\Phi$ is defined by
$\Phi^{n}(\alpha)=(\alpha|_{\mathcal{F}},\,(-1)^{n}\alpha|_{\mathcal{W}}).$
@@
where the alternating signs in $d_{\mathcal{F}}$ and $d_{\mathcal{W}}$ arise from incidence numbers, and the factor $(-1)^{n+1}$ arises solely from the degree shift in the identification $\Phi$.
@@ 5.1.4 Compatibility with Mapping Cones
The grounded diagnostics introduced earlier arise from applying a mapping cone construction to a grounding morphism $\varepsilon:\mathcal{F}\to\mathcal{W}$.
@@
$\mathcal{F}\overset{\varepsilon}{\rightarrow}\mathcal{W}\overset{\varepsilon^{\prime}}{\leftarrow}\mathcal{F}^{\prime}$
@@
It directly follows from the structure reduction (i.e., Corollary 1) and the induced long exact sequence of relative sheaf cohomologies (i.e., Lemma 1. ∎
@@ 5.2 Feasible Sets and Variational Characterization
Definition 10 ($\epsilon$-feasible set).
For $j\in\{0,1\}$ and $\epsilon\geq 0$, define
$\mathcal{S}_{j}(\epsilon)\;:=\;\big\{\,x\in C^{j}(K;\mathcal{F})\;:\;E_{j}(x)\leq\epsilon\,\big\},$
@@
$\mathcal{S}_{j}^{\mathrm{cone}}(\epsilon)\;:=\;\big\{\,x\;:\;E_{j}^{\mathrm{cone}}(x)\leq\epsilon\,\big\}.$
@@
For each $j\in\{0,1\}$, the families $\{\mathcal{S}_{j}(\epsilon)\}_{\epsilon\geq 0}$ and $\{\mathcal{S}_{j}^{\mathrm{cone}}(\epsilon)\}_{\epsilon\geq 0}$ form nested filtrations:
$0\leq\epsilon_{1}\leq\epsilon_{2}\quad\Longrightarrow\quad\mathcal{S}_{j}(\epsilon_{1})\subseteq\mathcal{S}_{j}(\epsilon_{2}),\qquad\mathcal{S}_{j}^{\mathrm{cone}}(\epsilon_{1})\subseteq\mathcal{S}_{j}^{\mathrm{cone}}(\epsilon_{2}).$
Interpretation. The parameter $\epsilon$ represents a budget of inconsistency. Increasing $\epsilon$ allows progressively larger violations of consistency or grounding constraints, yielding a persistence-style notion of feasibility.
Lemma 3 (Courant–Fischer Min–Max Theorem [1]).
Let $L_{j}$ have eigenvalues
$0=\lambda_{0}\leq\lambda_{1}\leq\lambda_{2}\leq\cdots.$
Then
$\lambda_{1}=\min_{\begin{subarray}{c}x\neq 0\\ x\perp\ker L_{j}\end{subarray}}\frac{E_{j}(x)}{\|x\|^{2}},$
and $\lambda_{k}$ is the smallest value of $\epsilon$ for which $\mathcal{S}_{j}(\epsilon)$ contains a $k$-dimensional linear subspace modulo $\ker L_{j}$. An analogous statement holds for $L_{j}^{\mathrm{cone}}$.
Proof.
Direct application of Min-Max theorem [1]. ∎
Interpretation. Eigenvalues represent the critical slack levels at which new feasible directions appear. Small eigenvalues correspond to weakly violated constraints, while large eigenvalues correspond to strongly violated constraints.
@@ 5.3.1 Global Spectral Witness
Rather than relying on a single eigenvalue, we aggregate information across the feasibility filtration.
@@
Fix $0\leq\epsilon_{0}<\epsilon_{1}$. For $j\in\{0,1\}$, define (global) spectral witness at degree $j$ by
$\mathcal{I}_{j}\;:=\;\int_{\epsilon_{0}}^{\epsilon_{1}}\Big(\min_{\begin{subarray}{c}x\neq 0\\ x\perp\ker L_{j}\end{subarray}}\frac{E_{j}(x)}{\|x\|^{2}}\;\wedge\;\epsilon\Big)\,d\epsilon,$
and analogously for the cone Laplacian, (global) coned spectral witness at degree $j$ is
$\mathcal{I}_{j}^{\mathrm{cone}}\;:=\;\int_{\epsilon_{0}}^{\epsilon_{1}}\Big(\min_{\begin{subarray}{c}x\neq 0\\ x\perp\ker L_{j}^{\mathrm{cone}}\end{subarray}}\frac{E_{j}^{\mathrm{cone}}(x)}{\|x\|^{2}}\;\wedge\;\epsilon\Big)\,d\epsilon,$
where $a\wedge b:=\min\{a,b\}$.
Interpretation. The quantity $\mathcal{I}_{j}$ measures how long infeasibility persists under increasing slack at degree $j$. The cone version $\mathcal{I}_{j}^{\mathrm{cone}}$ measures persistence of infeasibility once ambient grounding is enforced. Integrating over $\epsilon$ yields a persistence-style invariant without altering the base topology.
Remark 9 (Role of the kernel and spectral gaps).
The dimension of $\ker L_{j}$ encodes the exact cohomology group $H^{j}(K;\mathcal{F})$. Accordingly, all spectral witnesses are defined on the orthogonal complement of the kernel in order to factor out gauge symmetries associated with exact consistency. As a consequence, spectral witnesses measure deviation from exact consistency rather than consistency itself. In particular, a small spectral witness should not be interpreted as evidence of consistency, but rather as indicating that any inconsistency present is weak or poorly resolved at the given spectral scale.
Summary
The base Laplacians $(L_{0},L_{1})$ quantify ungrounded consistency and obstruction, while the cone Laplacians $(L_{0}^{\mathrm{cone}},L_{1}^{\mathrm{cone}})$ quantify grounded feasibility. The $\epsilon$-feasible sets induce a spectral filtration whose critical values are eigenvalues, and the resulting spectral witnesses provide a quantitative measure of counterfactual inconsistency.
@@ 5.3.2 Local spectral witnesses
Global spectral witnesses certify the existence and persistence of inconsistency, but they do not identify its spatial origin. To localize inconsistency to specific cells, we refine the $\epsilon$-feasible spectral filtration by decomposing consistency energy into cellwise contributions associated with low-energy spectral modes.
@@
Fix $\epsilon>0$ and define the spectral feasible subspace
$\mathcal{V}_{j}(\epsilon)\;:=\;\mathrm{span}\{\,v_{\ell}\;:\;\lambda_{\ell}\leq\epsilon\,\}\;\subseteq\;C^{j}(K;\mathcal{F}),$
@@
and let $w:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ be a nonnegative weighting function supported on a low-energy regime. For a spectral threshold $\epsilon>0$, we define the local spectral witness
$W_{j,\epsilon}:K_{j}\longrightarrow\mathbb{R}_{\geq 0}$
@@
$W_{j,\epsilon}(e):=\sum_{\lambda_{\ell}\leq\epsilon}w(\lambda_{\ell})\left(\sum_{\begin{subarray}{c}c_{j+1}\in K_{j+1}\\ c_{j+1}\geq e\end{subarray}}\bigl\|(d_{j}v_{\ell})[c_{j+1}]\bigr\|^{2}\;+\;\sum_{\begin{subarray}{c}c_{j-1}\in K_{j-1}\\ c_{j-1}\leq e\end{subarray}}\bigl\|(d_{j-1}^{T}v_{\ell})[c_{j-1}]\bigr\|^{2}\right).$
@@
An analogous definition applies to the cone Laplacian $L_{j}^{\mathrm{cone}}$, yielding $\mathcal{W}_{j,\epsilon}^{\mathrm{cone}}(c)$.
@@
  • For $j=0$, $\mathcal{W}_{0,\epsilon}(c)$ localizes inconsistency to edges (and, by aggregation, to nodes), measuring which local constraints are violated by low-energy modes.
  • For $j=1$, $\mathcal{W}_{1,\epsilon}(c)$ localizes obstruction both to (j+1)(j{+}1)-cells (triangle-level twists via $d_{1}$) and to edges (divergence or gauge components via $d_{0}^{\top}$).
  • Cone versions $\mathcal{W}_{j,\epsilon}^{\mathrm{cone}}$ localize inconsistency that persists under ambient grounding, isolating defects that are invisible to ungrounded analysis.
As $\epsilon$ increases, additional spectral modes contribute, yielding a persistence-style evolution of localized inconsistency. In this sense, $\mathcal{W}_{j,\epsilon}$ provides a cell-level refinement of the global spectral witness, identifying not only whether inconsistency exists, but where it is concentrated.
@@ 7.2 Magnitude: Degree of Inconsistency via Spectral Gap
When no global section exists up to spectral stability, we next ask how severe the inconsistency is. Table 3 compares a hidden twist and a noisy trivial bundle. Both have trivial kernel, but differ in their smallest nonzero eigenvalue and in the integrated spectral witness $\mathcal{I}_{0}$. These quantities measure the magnitude of inconsistency, i.e. the distance from admitting a global section, independent of spatial localization.
@@ 7.4 Relativity: Grounding-Induced Inconsistency via Mapping Cones
Figure 1: Local spectral witnesses for hidden twist (top row) and noisy trivial assignment (bottom row). Each row visualizes local spectral witnesses $\mathcal{W}_{j,\epsilon}$ on the same cycle complex, comparing intrinsic (base) Laplacian channels with grounded (mapping-cone) diagnostics. Shown are heatmaps induced by low-energy eigenmodes of the intrinsic $j=0$ and $j=1$ Laplacians, together with the relative cone channel corresponding to the intrinsic $j=1$ structure. Although the hidden twist and noisy trivial constructions exhibit comparable global spectral gaps, the hidden twist produces sharply localized low-energy modes, while the noisy trivial case displays diffuse energy. This localization behavior distinguishes structured logical defects from unstructured noise at the level of local spectral witnesses.

### v2

@@ 4.2.3 Equivalence under compatibility
Let $K$ be a finite simplicial complex and let $\epsilon:\mathcal{F}\to\mathcal{W}$ be a cellular sheaf morphism, by assuming the compatibility condition 4.2.1, and write $\epsilon^{\ast}:C^{\ast}(K;\mathcal{F})\to C^{\ast}(K;\mathcal{W})$ for the induced cochain map. Let $\widehat{K}$ denote the geometric cone on $K$ with apex vertex $\ast$, and let $\widehat{\mathcal{F}}$ be the cone sheaf on $\widehat{K}$ defined by \$4.2.2.
@@
$\Phi:C^{n}(\widehat{K};\widehat{\mathcal{F}})\cong\Bigl(\bigoplus_{\tau\in K_{n}}\mathcal{F}(\tau)\Bigr)\ \oplus\ \Bigl(\bigoplus_{\sigma\in K_{n-1}}\mathcal{W}(\sigma)\Bigr)\cong C^{n}(K;\mathcal{F})\ \oplus\ C^{n-1}(K;\mathcal{W})=\mathrm{Cone}(\epsilon^{\ast})[1]^{n}.$
@@
where the cone-cone contribution satisfies
$(d_{\widehat{\mathcal{F}}}\alpha)(\ast\tau)$ $=\sum_{\sigma\prec\tau}[\ast\tau:\ast\sigma]\,\mathcal{W}(\sigma\prec\tau)\,\alpha_{\mathcal{W}}(\ast\sigma)$
$=-\sum_{\sigma\prec\tau}[\tau:\sigma]\,\mathcal{W}(\sigma\prec\tau)\,y(\sigma)$
$=-(d_{\mathcal{W}}y)(\tau),\quad\forall\tau\in K_{n}.$
@@ 5.1.4 Compatibility with Mapping Cones
The grounded diagnostics introduced earlier arise from applying a mapping cone construction to a grounding morphism $\epsilon:\mathcal{F}\to\mathcal{W}$.
@@
$\mathcal{F}\overset{\epsilon}{\rightarrow}\mathcal{W}\overset{\epsilon^{\prime}}{\leftarrow}\mathcal{F}^{\prime}$
@@
It directly follows from the structure reduction (i.e., Corollary 1) and the induced long exact sequence of relative sheaf cohomologies (i.e., Lemma 1). ∎
@@ 5.2 Feasible Sets and Variational Characterization
Definition 10 ($\delta$-feasible set).
For $j\in\{0,1\}$ and $\delta\geq 0$, define
$\mathcal{S}_{j}(\delta)\;:=\;\big\{\,x\in C^{j}(K;\mathcal{F})\;:\;E_{j}(x)\leq\delta\,\big\},$
@@
$\mathcal{S}_{j}^{\mathrm{cone}}(\delta)\;:=\;\big\{\,x\;:\;E_{j}^{\mathrm{cone}}(x)\leq\delta\,\big\}.$
@@
For each $j\in\{0,1\}$, the families $\{\mathcal{S}_{j}(\delta)\}_{\delta\geq 0}$ and $\{\mathcal{S}_{j}^{\mathrm{cone}}(\delta)\}_{\delta\geq 0}$ form nested filtrations:
$0\leq\delta_{1}\leq\delta_{2}\quad\Longrightarrow\quad\mathcal{S}_{j}(\delta_{1})\subseteq\mathcal{S}_{j}(\delta_{2}),\qquad\mathcal{S}_{j}^{\mathrm{cone}}(\delta_{1})\subseteq\mathcal{S}_{j}^{\mathrm{cone}}(\delta_{2}).$
Interpretation. The parameter $\delta$ represents a budget of inconsistency. Increasing $\delta$ allows progressively larger violations of consistency or grounding constraints, yielding a persistence-style notion of feasibility.
@@ 5.3.1 Global Spectral Witness
To summarize the spectral filtration by scalar diagnostics, we introduce global spectral witnesses that aggregate low-energy spectral contributions across all feasible modes, quantifying the magnitude and persistence of inconsistency at a given cochain degree.
Let $\{\lambda_{\ell}^{(j)}\}_{\ell\geq 1}$ denote the nonzero eigenvalues of the $j$-th sheaf (or cone) Laplacian $L_{j}$, counted with multiplicity. Let $w:\mathbb{R}_{>0}\to\mathbb{R}_{\geq 0}$ be a fixed weighting function. Typical choices include
$w_{\mathrm{unif}}(\lambda)=1,\qquad w_{\mathrm{inv}}(\lambda)=\lambda^{-1},\qquad w_{\mathrm{heat}}(\lambda)=e^{-t\lambda}.$
@@
Fix $0\leq\delta_{0}<\delta_{1}$ and a nonnegative weighting function $w:\mathbb{R}_{>0}\to\mathbb{R}_{\geq 0}$. Let $\{\lambda^{(j)}_{\ell}\}_{\ell\geq 1}$ denote the nonzero eigenvalues of the $j$-th (regular or cone) Laplacian, counted with multiplicity. We define the multi-mode global spectral witness by
$\mathcal{I}_{j}(\delta_{0},\delta_{1})\;:=\;\sum_{\lambda^{(j)}_{\ell}\in(0,\delta_{1}]}(\delta_{1}-\max\{\delta_{0},\lambda^{(j)}_{\ell}\})\,w(\lambda^{(j)}_{\ell}).$
Special case: gap-based witness.
If the weighting function is chosen as $w(\lambda)=\mathbf{1}_{\{\lambda=\lambda_{\min}\}}$, then the global spectral witness reduces to a function of the smallest nonzero eigenvalue alone. This recovers the gap-based global witness used in earlier versions of this work, now interpreted as a single-mode special case of the general formulation.
In the diagnostic experiments reported in Section 7, this gap-based global spectral witness is used for simplicity and numerical stability; the general multi-mode definition is included to emphasize that the framework itself does not depend on privileging a single spectral mode.
Interpretation.
Each eigenmode contributes to the global spectral witness in proportion to the range of spectral slack over which it remains feasible. Modes with smaller eigenvalues persist longer as $\delta$ increases and therefore contribute more heavily to the witness. Under this formulation, the global spectral witness measures the total low-energy spectral mass associated with inconsistency, without privileging any single extremal eigenvalue.
@@ 5.3.2 Local spectral witnesses
Global spectral witnesses summarize how inconsistency accumulates across the spectral filtration, but they do not resolve where such inconsistency is supported spatially on the cell complex. To obtain cell-level diagnostics, we refine the same low-energy spectral structure by decomposing consistency energy across individual cells.
Local spectral witnesses assign to each cell a weighted contribution from low-energy spectral modes, yielding a spatial attribution of inconsistency that is consistent with the global spectral witness and evolves coherently with the spectral threshold.
@@
Fix $\delta>0$ and define the spectral feasible subspace
$\mathcal{V}_{j}(\delta)\;:=\;\mathrm{span}\{\,v_{\ell}\;:\;\lambda_{\ell}\leq\delta\,\}\;\subseteq\;C^{j}(K;\mathcal{F}),$
@@
and let $w:\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ be a nonnegative weighting function supported on a low-energy regime. For a spectral threshold $\delta>0$, we define the local spectral witness
$W_{j,\delta}:K_{j}\longrightarrow\mathbb{R}_{\geq 0}$
@@
$W_{j,\delta}(e):=\sum_{\begin{subarray}{c}\lambda_{\ell}\leq\delta\\ v_{\ell}\in\mathcal{V}_{j}(\delta)\end{subarray}}w(\lambda_{\ell})\left(\sum_{\begin{subarray}{c}e\prec c_{j+1}\in K_{j+1}\end{subarray}}\bigl\|(d_{j}v_{\ell})[c_{j+1}]\bigr\|^{2}\;+\;\sum_{\begin{subarray}{c}e\succ c_{j-1}\in K_{j-1}\end{subarray}}\bigl\|(d_{j-1}^{T}v_{\ell})[c_{j-1}]\bigr\|^{2}\right).$
@@
An analogous definition applies to the cone Laplacian $L_{j}^{\mathrm{cone}}$, yielding $\mathcal{W}_{j,\delta}^{\mathrm{cone}}(c)$.
@@
  • For $j=0$, $\mathcal{W}_{0,\delta}(c)$ localizes inconsistency to edges (and, by aggregation, to nodes), measuring which local constraints are violated by low-energy modes.
  • For $j=1$, $\mathcal{W}_{1,\delta}(c)$ localizes obstruction both to (j+1)(j{+}1)-cells (triangle-level twists via $d_{1}$) and to edges (divergence or gauge components via $d_{0}^{\top}$).
  • Cone versions $\mathcal{W}_{j,\delta}^{\mathrm{cone}}$ localize inconsistency that persists under ambient grounding, isolating defects that are invisible to ungrounded analysis.
As $\delta$ increases, additional spectral modes contribute, yielding a persistence-style evolution of localized inconsistency. In this sense, $\mathcal{W}_{j,\delta}$ provides a cell-level refinement of the global spectral witness, identifying not only whether inconsistency exists, but where it is concentrated.
@@ 7.2 Magnitude: Degree of Inconsistency via Spectral Gap
When no global section exists up to spectral stability, we next ask how severe the inconsistency is. Table 3 compares a hidden twist and a noisy trivial bundle. Both have trivial kernel, but differ in their smallest nonzero eigenvalue and in the integrated (gap-based) spectral witness $\mathcal{I}_{0}$. These quantities measure the magnitude of inconsistency, i.e. the distance from admitting a global section, independent of spatial localization.
@@ 7.4 Relativity: Grounding-Induced Inconsistency via Mapping Cones
Figure 1: Local spectral witnesses for hidden twist (top row) and noisy trivial assignment (bottom row). Each row visualizes local spectral witnesses $\mathcal{W}_{j,\delta}$ on the same cycle complex, comparing intrinsic (base) Laplacian channels with grounded (mapping-cone) diagnostics. Shown are heatmaps induced by low-energy eigenmodes of the intrinsic $j=0$ and $j=1$ Laplacians, together with the relative cone channel corresponding to the intrinsic $j=1$ structure. Although the hidden twist and noisy trivial constructions exhibit comparable global spectral gaps, the hidden twist produces sharply localized low-energy modes, while the noisy trivial case displays diffuse energy. This localization behavior distinguishes structured logical defects from unstructured noise at the level of local spectral witnesses.
@@ B.1 Input Data and Ambient Representation
Remark 10.
Although node-level features are used throughout the experiments in this paper, the framework does not assume that the hypothetical structure is vertex-generated.
@@ B.7 Spectral Witnesses
B.8 Hypothetical structures and stability across admissible sheaf realizations
We formalize the sense in which the proposed spectral diagnostics are independent of specific sheaf-construction choices, while remaining sensitive to the underlying structural hypotheses.
Hypothetical structure.
A hypothetical structure $\mathcal{H}$ consists of partial cellular data specified on a cell complex $K$, including:
  • a choice of cells $\Sigma\subseteq K$ on which stalk data are prescribed;
  • for each $\sigma\in\Sigma$, a finite-dimensional inner product space $\mathcal{H}(\sigma)$ representing structural features at $\sigma$;
  • optionally, partial compatibility data or constraints between prescribed stalks.
No assumption is made that $\Sigma$ consists only of vertices, nor that restriction maps are fully specified. In particular, $\mathcal{H}$ need not define a sheaf.
Family of constructions.
Given a hypothetical structure $\mathcal{H}$, we consider a family of admissible construction procedures indexed by parameters $\theta\in\Theta$. Each construction specifies:
  • how to extend stalk assignments from $\Sigma$ to additional cells when required;
  • how to choose bases or coordinate representations for stalks;
  • how to deduce or approximate restriction maps between cells;
  • how to select auxiliary parameters such as tolerances or truncation thresholds.
These choices are algorithmic and modeling-dependent, and are not regarded as intrinsic to $\mathcal{H}$.
Realized sheaves and admissible objects.
Each construction $\theta\in\Theta$ produces a cellular sheaf
$\mathcal{F}_{\theta}\;=\;\mathcal{C}_{\theta}(\mathcal{H})$
on $K$, equipped with inner product stalks and linear restriction maps. Different choices of $\theta$ may yield non-isomorphic sheaves, supported on different collections of cells or with different internal dimensions.
We denote by $\mathsf{Shv}(\mathcal{H})$ the resulting class of admissible realizations of $\mathcal{H}$, i.e. the collection of cellular sheaves arising from constructions compatible with the same hypothetical structure. No canonical morphisms between elements of $\mathsf{Shv}(\mathcal{H})$ are assumed to exist.
Spectral indicators.
Let $\mathcal{D}$ denote any spectral diagnostic derived from the sheaf Laplacians (or cone Laplacians), such as spectral gaps, integrated spectral witnesses, or localized energy profiles. The framework does not assert that $\mathcal{D}(\mathcal{F}_{\theta})$ is invariant under changes of $\theta$. Rather, it asserts that for admissible constructions $\theta,\theta^{\prime}\in\Theta$, the quantities
$\mathcal{D}(\mathcal{F}_{\theta})\quad\text{and}\quad\mathcal{D}(\mathcal{F}_{\theta^{\prime}})$
are quantitatively comparable and stable up to explicitly controlled scale parameters.
Interpretation.
In this sense, the spectral diagnostics are understood as properties of the underlying hypothetical structure $\mathcal{H}$, as revealed through a class of admissible sheaf realizations, rather than invariants of any single constructed sheaf. We refer to this principle as stability across admissible sheaf realizations.